Pavan, Nate, Eric, Dominick, Gavin, Craig, Chris, Ryan

FTS (Nate)

Nate wrote a script for restarting FTS — makes life easier for restarting all three FTS servers
Anyone can do it (with novapro priviliges) — just send a courtesy email to the group.

Data handling requests

ifdh_fetch behaving differently from different packages — writing single file to app area
/tmp isn’t big enough — we need to follow up.
More info in an email…

ND Cosmics (Eric)

— generation status:
- 1st time — copy out failed
- 2nd time — Another failed copy out…
- 3rd time — duplicate parents
- 4th time - s14–03-24, and correct incantation for DCACHE, but no hash directories.
—> Nate ran script to rename to proper name and restarted FTS for pushing files to SAM
—> FD used 14-02-05
- 223 cosmic files — ~2000 files is the full dataset
- Using FTS 01 for these —> issue with “Poisson lambda changes”
- log files aren’t available due to condor_tmp filling up
- S14-05-12 will be used — will do all ND with this tag.

PCHits status (Satish)

Satish will start tests today.
Initial tests ran into a multipoint snag. This should go away once ND cosmics are processed in S14-05-12.

ND Overlay (Nate/Eric)

Can’t do reconstruction of this until we have ND cosmics.
Some issues with running offsite — some libraries missing offsite? Gavin looking into this.
Gavin will run ND overlays after backports in S14-05-12 so all ND generation is in the same tag.

FD plan (Gavin/Ryan)

Some misleading information from analysis groups.
— Reprocess 4DB for nu_mu group
— nu_e group would like 7DB data (10x what they used for neutrino)
( Only similar amount of livetime for contiguous 7-DB data)
— Instead could increase 4 DB MC.
— Still outstanding : work with Reco to make sure the latest software suite has the fixes.
— Can we use the same tag as the 4DB cosmic sample for the 4DB cosmic data.
— 7 DB calibration is not running now due to memory issues. No ETA.
— Nate will double check what tag we used for the 4 DB MC…
Dominick will run reco on more 4DB data and eventually 7DB data. Previously only ran 10 subruns so can increase to pick
up all 64 subruns. Use tag S14-05-08. PID in S14-05-12.